Circular No. 8899

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 W1 (BOATTINI) A. Boattini reports the discovery of a comet found on CCD images

taken with the Mt. Lemmon Survey’s 1.5-m reflector (discovery observation tabulated below), noting that four co-added 30-s exposures show a slightly asymmetric condensation of size ∼ 8′′ × 6′′ with the longer axis in p.a. 290^{o}–300^{o}. Boattini adds that ten co-added 90-s CCD exposures taken by C. Snodgrass with the European Southern Observatory’s 3.58-m New Technology Telescope (+ EMMI) on Nov. 23.34–23.36 UT show a coma of size ∼ 6′′ × 7′′ and a very faint tail in p.a. 293^{o} extending ∼ 9′′ from the comet’s center. Following posting on the Minor Planet Center’s NEOCP webpage, E. Guido and G. Sostero (Castellammare di Stabia, Italy) report that CCD images obtained remotely with a 0.25-m f/3.4 reflector near Mayhill, NM, U.S.A., on Nov. 21.5 shows (via 30 co-added exposures) a coma with diameter ∼ 12′′ that is elongated toward the southeast.

2007 UT  α_{2000}  δ_{2000}  Mag.
Nov. 20.48331  11^{h}45^{m}54.^{s}46  +0^{o}18′59′′.7  18.1

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-W63.

T = 2008 June 14.300 TT  ω = 310.^{o}982
Ω = 334.585 }2000.0
q = 0.72012 AU  i = 10.352

V598 PUPPIS G. Pojmanski, D. Szczygiel, and B. Pilecki, Warsaw University As-

tronomical Observatory, report the following prediscovery V magnitudes for this nova (cf. IAUC 8898) from 3-min CCD exposures obtained with a 70-mm (200-mm-f.l.) f/2.8 telephoto lens in the course of the All Sky Automated Survey (pixel size 14′′.8), indicating that the object erupted in early June and reached V ∼ 4.1 (when the images were saturated, and aperture photometry often underestimates the true stellar brightness by 0.1-0.5 mag): May 8.059 UT, [14:; 12.996, [14:; June 1.966, [14:; 5.968, 4.08; 17.954, 6.50; Aug. 18.429, 9.28; 22.412, 9.33; 28.396, 9.58; Sept. 1.415, 9.44; 5.408, 9.49; 11.399, 9.56; 19.383, 9.69; 22.390, 9.74; 26.355, 9.77; 30.358, 9.86; Oct. 4.350, 9.93; 8.343, 9.98; 12.329, 10.02; Nov. 4.281, 10.36; 7.329, 10.43; 14.353, 10.45; 18.261, 10.55. The nova also was not visible in an I-band image taken on June 2.978.

2007 November 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green